# Fractional coloring via entropy

### 18 Mar 2026 v1

@@ 1.1. Background and main results
$\chi_{f}(H)\leq\left(1+\varepsilon\right)\left(\frac{r}{r-1}\right)\left(r(r-1)\dfrac{d}{\log d}\right)^{\frac{1}{r-1}}.$
@@
$\mathbb{P}[v\in I]\geq\left(1-\varepsilon\right)\left(1-\frac{1}{r}\right)\left(\dfrac{\log d}{r(r-1)d}\right)^{\frac{1}{r-1}},\qquad\text{for each }v\in V(H).$
@@
$\chi_{f}(H)\leq\left(1+\varepsilon\right)\left(\frac{r}{r-1}\right)\left(\frac{r(r-1)(3r-4)}{(3r-6)}\dfrac{d}{\log d}\right)^{\frac{1}{r-1}}.$
@@ 1.3. Proof Overview
The algorithm takes as input a $d$-degenerate (hyper)graph $H$ and a parameter $q\in\mathbb{N}$ and outputs a collection of subsets $\left\{S(v)\subseteq[q]\right\}_{v\in V(H)}$ such that $\cap_{u\in e}S(u)=\emptyset$ for each edge $e\in E(H)$. We initially assign each vertex-color pair $(v,c)$ a weight $p_{0}(v,c)=\alpha$ for a well-chosen $\alpha$ (our eventual bound on $\chi_{f}(H)$ is $\Theta(\alpha^{-1})$). Iterating through the vertices in a degeneracy ordering of $H$, we determine the set $S(v_{i})$ by selecting $c\in[q]$ with a probability determined by $p_{i-1}(v_{i},c)$ (we are being intentionally vague at this point as there is a distinction in how we do so for each of our results). We update the weights $p_{i}(v_{k},c)$ for each $k>i$ and $c\in[q]$ according to the outcome of the event $\left\{c\in S(v_{i})\right\}$. In particular, the weight $p_{i}(v_{k},c)$ must be set to 0 if some edge $e\ni v_{k}$ satisfies $c\in\cap_{u\in e\setminus\{v_{k}\}}S(u)$. Additionally, we increase the weight for $v_{k}\in N_{R}(v_{i})$ if $c\notin S(v_{i})$ as we would like to make it more likely that $c\in S(v_{k})$ in this case. Our choice of parameters ensures that $\mathbb{E}[p_{i}(v_{k},c)\mid p_{i-1}(\cdot,\cdot)]=p_{i-1}(v_{k},c)$. It then follows that $\mathbb{E}[|S(v_{k})|]=q\alpha$.
@@ 1.4. Open problems
Here, $\log^{(b)}(\cdot)$ denotes the $b$-fold iterated logarithm function defined as $\log^{(b)}(x)\coloneqq\log(\log^{(b-1)}(x))$ and $\log^{(1)}(x)=\log x$. They ask whether $d$-degenerate $K_{r}$-free graphs $G$ satisfy $\chi_{f}(G)=o_{r}(d)$ for $r\geq 4$ in general [11, Problem 6.4]. It is possible that the techniques described in this paper could yield an affirmative answer to this question.
@@ 2.1. Algorithm description
$\chi_{f}(H)\leq\left(1+\varepsilon\right)\left(\frac{r}{r-1}\right)\left(r(r-1)\dfrac{d}{\log d}\right)^{\frac{1}{r-1}}.$
@@
$\mathbb{P}[v\in I]\geq\left(1-\varepsilon\right)\left(1-\frac{1}{r}\right)\left(\dfrac{\log d}{r(r-1)d}\right)^{\frac{1}{r-1}},\qquad\text{for each }v\in V(H).$
@@ 3.1. Algorithm description
    Initialize: For each $v\in V(H)$, set $B_{0}^{\mu}(v)=B_{0}^{\ell}(v)=\emptyset$. For each $v\in V(H)$ and $c\in[q]$, set $p_{0}(v,c)=\alpha$, where $\alpha\coloneqq\left(\frac{\log d}{(1+\varepsilon r/10)r(r-1)d}\right)^{\frac{1}{r-1}}$. Additionally, fix a degeneracy ordering $(v_{1},\ldots,v_{n})$ of $V(H)$.
@@
$\mathbb{P}\left[\exists k\text{ s.t. }|S(v_{k})|\leq\frac{q}{(1+\varepsilon)}\left(1-\frac{1}{r}\right)\left(\dfrac{\log d}{r(r-1)d}\right)^{\frac{1}{r-1}}\right]$ $\leq\mathbb{P}\left[\exists k\text{ s.t. }|S(v_{k})|\leq(1-\varepsilon/20)\left(1-\frac{1}{r}\right)q\alpha\right]$
@@
$\frac{(r-1)d\alpha^{r-1}}{\log\left(1/\left(2\alpha^{1+\kappa}\right)\right)}\leq\frac{\log d}{r(1+\varepsilon\,r/10)(1+\kappa)\log\left(\frac{r(r-1)(1+\varepsilon r/10)d}{2\log d}\right)}\leq\frac{1}{r}.$

### 13 Apr 2026 v2

@@ 1.1. Background and main results
$\chi_{f}(H)\leq\left(1+\varepsilon\right)\left(\frac{r}{r-1}\right)\left(r(r-1)^{2}\dfrac{d}{\log d}\right)^{\frac{1}{r-1}}.$
@@
$\mathbb{P}[v\in I]\geq\left(1-\varepsilon\right)\left(1-\frac{1}{r}\right)\left(\dfrac{\log d}{r(r-1)^{2}d}\right)^{\frac{1}{r-1}},\qquad\text{for each }v\in V(H).$
@@
$\chi_{f}(H)\leq\left(1+\varepsilon\right)\left(\frac{r}{r-1}\right)\left(\frac{r(r-1)^{2}(3r-4)}{(3r-6)}\dfrac{d}{\log d}\right)^{\frac{1}{r-1}}.$
@@ 1.3. Proof Overview
The algorithm takes as input a $d$-degenerate (hyper)graph $H$ and a parameter $q\in\mathbb{N}$ and outputs a collection of subsets $\left\{S(v)\subseteq[q]\right\}_{v\in V(H)}$ such that $\cap_{u\in e}S(u)=\emptyset$ for each edge $e\in E(H)$. We initially assign each vertex-color pair $(v,c)$ a weight $p_{0}(v,c)=\alpha$ for a well-chosen $\alpha\in[0,1]$ (our eventual bound on $\chi_{f}(H)$ is $\Theta(\alpha^{-1})$). Iterating through the vertices in a degeneracy ordering of $H$, we determine the set $S(v_{i})$ by selecting $c\in[q]$ with a probability determined by $p_{i-1}(v_{i},c)$ (we are being intentionally vague at this point as there is a distinction in how we do so for each of our results). We update the weights $p_{i}(v_{k},c)$ for each $k>i$ and $c\in[q]$ according to the outcome of the event $\left\{c\in S(v_{i})\right\}$. In particular, the weight $p_{i}(v_{k},c)$ must be set to 0 if some edge $e\ni v_{k}$ satisfies $c\in\cap_{u\in e\setminus\{v_{k}\}}S(u)$. Additionally, we increase the weight for $v_{k}\in N_{R}(v_{i})$ if $c\notin S(v_{i})$ as we would like to make it more likely that $c\in S(v_{k})$ in this case. Our choice of parameters ensures that $\mathbb{E}[p_{i}(v_{k},c)\mid p_{i-1}(\cdot,\cdot)]=p_{i-1}(v_{k},c)$. It then follows that $\mathbb{E}[|S(v_{k})|]=q\alpha$.
@@ 1.4. Open problems
$\chi_{f}(G)\geq c_{r}\frac{d}{\log^{(r-2)}d}$ for some constant $c_{r}>0$ [11, Theorem 1.6] Here, $\log^{(b)}(\cdot)$ denotes the $b$-fold iterated logarithm function defined as $\log^{(b)}(x)\coloneqq\log(\log^{(b-1)}(x))$ and $\log^{(1)}(x)=\log x$. They ask whether $d$-degenerate $K_{r}$-free graphs $G$ satisfy $\chi_{f}(G)=o_{r}(d)$ for $r\geq 4$ in general [11, Problem 6.4]. It is possible that the techniques described in this paper could yield an affirmative answer to this question.
@@ 2.1. Algorithm description
$\chi_{f}(H)\leq\left(1+\varepsilon\right)\left(\frac{r}{r-1}\right)\left(r(r-1)^{2}\dfrac{d}{\log d}\right)^{\frac{1}{r-1}}.$
@@
$\mathbb{P}[v\in I]\geq\left(1-\varepsilon\right)\left(1-\frac{1}{r}\right)\left(\dfrac{\log d}{r(r-1)^{2}d}\right)^{\frac{1}{r-1}},\qquad\text{for each }v\in V(H).$
@@ 3.1. Algorithm description
    Initialize: For each $v\in V(H)$, set $B_{0}^{\mu}(v)=B_{0}^{\ell}(v)=\emptyset$. For each $v\in V(H)$ and $c\in[q]$, set $p_{0}(v,c)=\alpha$, where $\alpha\coloneqq\left(\frac{\log d}{(1+\varepsilon r/10)r(r-1)^{2}d}\right)^{\frac{1}{r-1}}$. Additionally, fix a degeneracy ordering $(v_{1},\ldots,v_{n})$ of $V(H)$.
@@
$\mathbb{P}\left[\exists k\text{ s.t. }|S(v_{k})|\leq\frac{q}{(1+\varepsilon)}\left(1-\frac{1}{r}\right)\left(\dfrac{\log d}{r(r-1)^{2}d}\right)^{\frac{1}{r-1}}\right]$ $\leq\mathbb{P}\left[\exists k\text{ s.t. }|S(v_{k})|\leq(1-\varepsilon/20)\left(1-\frac{1}{r}\right)q\alpha\right]$
@@
$\frac{(r-1)d\alpha^{r-1}}{\log\left(1/\left(2\alpha^{1+\kappa}\right)\right)}\leq\frac{\log d}{r(1+\varepsilon\,r/10)(1+\kappa)\log\left(\frac{r(r-1)^{2}(1+\varepsilon r/10)d}{2\log d}\right)}\leq\frac{1}{r}.$